y(0) = 2, 70) = 2 (1 point) Use the Laplace transform to solve the following initial value problem: y – 11y + 30y = 0, (1)
Answer #1

Soin yll-11y+30y=0 y (o)=2 or y co)=2 we know that Now yll- 119 +30y zo where takking Laplace transform both side şs²y-Sy(Solving the partial fraction form 2S- 20 B (5-6) (8-5) (56) earis 25-20 (5-6) (5-5) ACS-5) +B (5-6) (86) (8-5) eq ci puttingLa place Inverse Now takking Işyesi ? 8-1-f= + JH 10 (9-3) LS-5) = -8e 79 (+) B + 10 estif you have any doubt regarding this please let me know and if you understand the solution than give me a thumbs up thanks